### v1

@@ Abstract
Differentially private stochastic gradient descent (DP-SGD) is the standard algorithm for training machine learning models under differential privacy (DP). The major drawback of DP-SGD is the drop in utility which prior work has comprehensively studied. However, in practice another major drawback that hinders the large-scale deployment it is the significantly higher computational cost. We conduct a comprehensive empirical study to quantify the computational cost of training deep learning models under DP and benchmark methods that aim at reducing the cost. Among these are more efficient implementations of DP-SGD and training with lower precision. Finally, we study the scaling behaviour using up to 80 GPUs.
@@ 1 Introduction
Training data of machine learning (ML) models can be extracted Balle et al. (2022); Carlini et al. (2021). Differential Privacy (DP) (Dwork et al., 2006) is the gold standard for formalizing the privacy leakage of training data in ML and mitigating the risk of privacy attacks on the training data. DP is deployed in many applications involving sensitive data Abowd (2018); Cormode et al. (2018).
The established algorithm for integrating DP into the training pipeline of deep learning models is DP stochastic gradient descent (DP-SGD) Rajkumar & Agarwal (2012); Song et al. (2013); Abadi et al. (2016), which is the DP adaptation of SGD (see also Algorithm 1). DP-SGD has two major drawbacks in comparison to SGD:
Higher computational cost: DP-SGD requires more memory and is computationally more expensive due to the per-sample clipping.
@@
We find that non-optimized training with DP-SGD costs up to three and eight times more than non-private training for ViT and ResNets, respectively (See Section 4).
@@
We identify the reasons that lead to the higher computational cost of DP-SGD using profiling (See Section 4.3).
@@
We benchmark different strategies that can reduce this cost drastically up to a level that matches the non-optimized non-private training:
More efficient implementations of DP-SGD (See Section 5.1).
Use of JAX instead of PyTorch (See Section 5.2).
Lower Precision with TF-32 (See Section 5.3).
@@ 2.1 DP-SGD Algorithm
Poisson subsampling Interestingly, Bu et al. (2022) and Bu et al. (2023) never mention Poisson subsampling in their works of Mix Ghost Clipping and Book Keeping. Even more, Bu et al. (2022) states that it has a speed-up of times against other algorithms with a fixed batch size, which would affect the privacy accountant method. The same happens in practice for JAX implementations De et al. (2022), were the sampling is done by shuffling the dataset and using each sample once per epoch. While it is easier to implement in practice, it does not use the correct Poisson subsampling for the numerical accounting methods.
To make a fair comparison between all methods, we implement the Poisson subsampling, the same way Opacus does it, for all frameworks and a new custom Batch Memory Manager to flag when it is time to take a step. This way, all the experiments are seeded and will have the same logical and physical batch sizes.
2.2 DP-SGD Optimizations
@@
JAX We compare all implementations with a native JAX Bradbury et al. (2018) implementation that clips the per-sample gradients with Optax DeepMind et al. (2020) without utilizing any further optimization.
Ghost clipping computes the gradient norm loss after the backpropagation optimization and then reweights the loss to update the clipped gradients. Its main contribution is memory saving at the cost of adding another backward pass (Li et al., 2022).
Mixed Ghost clipping proposed by Bu et al. (2022) builds on-top of Ghost clipping. It implements the ghost clipping technique for convolutional layers. Its main contribution is that the algorithm will decide when to clip the gradients using per-example or ghost. This difference matters because the ghost clipping is less efficient when the layer’s input dimensions are too big. E.g., for ResNets, each clipping method will be applied for half of the layers. The first layers will be clipped using the the per-example and then ghost clipping in the bottom layers. As the model goes deeper, the feature size decreases, and the number of channels increases, prioritizing ghost clipping (Bu et al., 2022).
Book keeping by Bu et al. (2023) uses all the previous techniques but requires only one backpropagation pass without explicitly calculating the per-example gradients. It avoids the second pass that ghost clipping does by reusing the intermediate results of the output gradients to calculate the sum of the clipped gradients and the clipping factor. Book keeping can also be implemented together with the mix optimization. It also implements another technique called MixOpt, which does the same evaluation as the mix ghost clipping, but also determines whether doing a second backward pass is more efficient.
@@ 3 Experiment Overview
In our experiments we compare the throughput, defined as how many samples can be processed per second during training.
Datasets We benchmark with the dataset CIFAR100 (Krizhevsky & Hinton, 2009). It contains 60,000 data points, from 100 classes. It is a commonly used dataset for testing computer vision models and their private counterparts (Abadi et al., 2016; De et al., 2022; Yousefpour et al., 2021).
Models We benchmark two families of models. The Vision Transformer (ViT) (Dosovitskiy et al., 2021) and BiT-ResNet (Kolesnikov et al., 2020) (SeeTable 1). All models are pre-trained on the same public dataset, ImageNet-21k. However, the ViT models are additionally fine-tuned on ImageNet-1k (Steiner et al., 2022; Wightman, 2019). For our experiments, we are fine-tuning the feature encoder on CIFAR100. CIFAR-100 images are resized to 224x224.
Parameterization While parameter-efficient fine-tuning of some parts of the model, e.g. using adapters like FiLM Perez et al. (2018) or LoRA (Hu et al., 2022) has been shown to be effective under DP Tobaben et al. (2023); Yu et al. (2022), our work focuses on the computational efficiency of DP-SGD and thus we consider on the worst-case scenario which is fine-tuning all parameters of the model. Furthermore, any training from scratch requires training all parameters.
| Model | # of Parameters | |
| ViT | Tiny | 5.7 M |
| Small | 22.1 M | |
| Base | 86.6 M | |
| Large | 304.3 M | |
| Huge | 630.8 M | |
| BiT-ResNet | 23.7 M | |
| 42.7 M | ||
| 211.8 M | ||
| 382.4 M | ||
| 929.2 M | ||
Hyperparameters All experiments use the optimal hyperparameters found previously by Tobaben et al. (2023) with a privacy budget of . We train for two epochs, and according to the sampling rate and batch sizes, there will be four optimizer steps in total. Using only two epochs allows us to test the experiments quickly and evaluate them over multiple runs. We do not focus on finding the best possible utility, which requires training for many more epochs. In Table A4 we show the mean accuracy after training for just two epochs. For optimal utility, it must be trained for a higher number of epochs.
Environment specifications We use two GPU architectures, NVIDIA V100 and NVIDIA A100, with 32 and 40 GB of VRAM respectively. Both use identical Python environments, duplicated in both to make a fair comparison. For the multinode experiments, each node has four GPU units. All experiments use 16 CPU workers, except the distributed DP, which breaks with more than one worker.
Source code We will make the source code for replicating the experiments available at https://github.com/DPBayes/Towards-Efficient-Scalable-Training-DP-DL.
@@ 4 What is the cost of DP in Deep Learning?
In this section we will quantify the computational cost of deploying DP training as a whole. We do this by comparing the maximum physical batch sizes and throughputs between the non-private training and Opacus per-example clipping, called O-flat, which is the most used DP-SGD implementation. Additionally, we identify the reasons for the higher computational cost of DP-SGD through profiling.
4.1 Maximum Physical Batch Size
We first compare how many samples we can process in one physical batch before running out of memory. In Figure 2, we can see that the model size matters as its parameters must be stored in memory, and their gradients must also be stored for accumulation. As there are millions of parameters, it becomes increasingly expensive to store them in memory. As expected, due to its implementation of calculating the gradients for each sample, the per-example clipping fits fewer instances.
The difference in throughput between non-private and private training grows with respect to the model size. In Figure 2(a), the difference for the huge model size is already 11x times the number of fitted samples, and the private model can only fit three samples, getting closer to just calculating one sample at a time and accumulating that for a large logical batch, which is very inefficient. This is problematic as a bigger model would be unable to fit even one sample.
From the experiments we also verified that the use of physical batches does not affect the model performance; each model size will reach a different accuracy than the others, but the same model size will always reach the same performance independent of the physical batch size.
4.2 Throughput Comparision
The batch size alone does not tell the whole story, as the main interest is how much slower the private training is. Figure 3 shows the relative difference between mean throughput, which is how faster the non-private training is and how much more expensive the private one is. It is consistent with Figure 2 in that as the model size grows, the difference between how many samples can be fitted and the throughput difference will grow larger.
@@ 4.2 Throughput Comparision
ViT The throughput difference between Opacus and a non-private baseline for the ViT model (see Figure 3(a)) does not spike and grows steadily, which is interesting considering how big the relative difference is in physical batch size. Using the previous example with the ViT huge model, the throughput difference is about , while we can fit 11 times more samples in the non-private implementation.
@@
ResNets On the other hand, in Figure 3(b), the ResNet models have different behavior. They do have spikes of growth as the model size grows. The contrast in Figure 3 between ViT and BiT ResNet models is due to the architecture and types of layers. The parameter space grows as the width factor (see Table 1) for the ResNets, so the makes the neural network wider by a factor of three. Based on our results, the width of the layers affects throughput much more than the depth of the network. They have comparable throughput with the same width and different depths, but increasing the width will make the model in the private setting much slower and reduce the maximum batch size significantly.
How much does finding the maximum physical batch size matter? In Figure 4, we display the relative throughput as a percentage by dividing the throughput at a particular physical batch size by the maximum achievable throughput. We see that as the physical batch size increases, the throughput will grow as expected, but at some point there is no signifant further improvement in throughput from using a larger physical batch size.
@@
Practitioners may estimate the optimal batch size based on available memory and performance trade-offs. It is not crutial to set the physical batch size to the maximum possible but a good enough value be fine. Typically, smaller batches are limited by data loading speeds, but as batch size increases, computation becomes the limiting factor.
@@ 4.3 Reasons for the Increase in Computational Cost
Giving a detailed breakdown of low-level operations associated with DP is challenging. However, using GPU profiling tool NVIDIA Nsight System we can identify three aspects which constitute to the majority DP overheads. Firstly, due to its larger memory footprint, DP-SGD is able to consume smaller physical batches than its non-private counterpart. This results in larger amount of smaller low-level kernel calls which leads to slightly lower utilisation of the GPU compute. At very small batch sizes even the kernel launch overheads can become a notable factor for slowdown. Secondly, the computation of per-example-gradients introduces significant overhead in the backward pass as it cannot be parallelised as in batched gradient computation. This is the most prominent cause of the total overhead. Finally, an additional DP-optimizer step that clips and accumulates the per example gradients, which is not present in the non-dp algorithm, needs to taken after each physical batch (see Table A3).
@@ 5 Decreasing the computational cost
5.1 Optimized algorithms
First, we evaluate the more efficient implementations that have been described in Section 2.2 using the ViT base model. We chose it as our benchmark model because the middle model size is large enough to test our hypothesis but less extensive than it would take too much time to train. The implementations do not support the BiT ResNet due to their custom weight standardization layer.
@@ 5.1 Optimized algorithms
Throughput Comparision Figure 5 displays the throughput for each clipping algorithm for each available GPU. To understand it better, we also compute the relative difference between the mean throughput. By using better hardware components, in this case, improving from a V100 to an A100 GPU has an increment of times in throughput for all clipping methods. The one that benefits the most is the per-example clipping by Opacus with a improvement in throughput. It is because of Opacus-specific optimizations. Their implementation is optimized to vectorize the virtual batches and get the most out of the processing unit (Yousefpour et al., 2021) to compensate for the per-example clipping. Also, we base our virtual batching module on Opacus, being specially optimized for the Opacus implementation, yielding better results with better hardware. The other implementations will have benefits similar to those of non-private training. The clipping methods in both GPU architectures will have the same relative throughput regarding the non-private performance. Private Vision gets closer to the non-private baseline physical batch size, but Book Keeping is closer to its throughput with a smaller physical batch size (see Figure 6).
@@
5.2 JAX framework

### v2

@@ Abstract
Differentially private stochastic gradient descent (DP-SGD) is the standard algorithm for training machine learning models under differential privacy (DP). The most common DP-SGD privacy accountants rely on Poisson subsampling for ensuring the theoretical DP guarantees. Implementing computationally efficient DP-SGD with Poisson subsampling is not trivial, which leads to many implementations ignoring this requirement. We conduct a comprehensive empirical study to quantify the computational cost of training deep learning models under DP given the requirement of Poisson subsampling, by re-implementing efficient methods using Poisson subsampling and benchmarking them. We find that using the naive implementation DP-SGD with Opacus in PyTorch has between 2.6 and 8 times lower throughput of processed training examples per second than SGD. However, efficient gradient clipping implementations with e.g. Ghost Clipping can roughly halve this cost. We propose alternative computationally efficient ways of implementing DP-SGD with JAX that are using Poisson subsampling and achieve only around 1.2 times lower throughput than SGD based on PyTorch. We highlight important implementation considerations with JAX. Finally, we study the scaling behaviour using up to 80 GPUs and find that DP-SGD scales better than SGD. We share our re-implementations using Poisson subsampling at https://github.com/DPBayes/Towards-Efficient-Scalable-Training-DP-DL.
@@ 1 Introduction
Training data of machine learning (ML) models can be vulnerable to extraction (Balle et al., 2022; Carlini et al., 2021). Differential Privacy (DP) (Dwork et al., 2006) is the gold standard for formalizing the privacy leakage of training data in ML and mitigating the risk of privacy attacks on the training data. DP is deployed in many applications involving sensitive data (Abowd, 2018; Cormode et al., 2018).
The established algorithm for integrating DP into the training pipeline of deep learning models is DP stochastic gradient descent (DP-SGD) (Rajkumar & Agarwal, 2012; Song et al., 2013; Abadi et al., 2016), which is the DP adaptation of SGD (see also Algorithm 1). DP-SGD has two major drawbacks in comparison to SGD: higher computational cost and loss in utility. DP-SGD requires more memory and is computationally more expensive due to the per-example clipping. The utility in comparison to non-DP training drops but this can be mitigated to certain extend by using larger batch sizes (Räisä et al., 2024) and training longer (Ponomareva et al., 2023) which further increases the computational cost.
Standard DP bounds assume so-called Poisson subsampling, where each example is selected at each iteration independently with a fixed probability. This implies that different minibatches will be of different size, and makes efficient implementation more difficult. As a result, many existing implementations forego proper implementation of Poisson subsampling. Recent research (Lebeda et al., 2024) shows that such implementations may have significantly weaker privacy guarantees than claimed under Poisson subsampling assumption.
@@
We re-implement all DP-SGD methods with Poisson subsampling that is fully DP and share the source code at https://github.com/DPBayes/Towards-Efficient-Scalable-Training-DP-DL
@@
We find that non-optimized training with DP-SGD costs betweem 2.6 and 3.2 times more than non-private training for ViT and 4 to 8 times for ResNets (See Section 4). We identify the reasons that lead to the higher computational cost of DP-SGD using profiling.
@@
We benchmark different strategies that can reduce this cost drastically up to a level that matches the non-optimized non-private training (See Figure 1 for an overview): 1 More efficient gradient clipping implementations of DP-SGD (See Section 5.1). 2 Lower Precision with TF32 (See Section 5.2).
@@
We propose a JAX implementation relying on proper Poisson sampling that is not prone to re-compilation and outperforms a naive implementation in terms of throughput (See Section 6).
5.
@@ 2.1 DP-SGD Algorithm
Poisson subsampling Interestingly, Bu et al. (2022) and Bu et al. (2023) never mention Poisson subsampling in their works of Mix Ghost clipping and Book Keeping. Even more, Bu et al. (2022) state that it has a speed-up of times against other algorithms with a fixed batch size, which would affect the privacy accounting method. The same happens in practice for JAX implementations (De et al., 2022), where the sampling is done by shuffling the dataset and using each sample once per epoch. While this makes efficient implementation easier, it does not use the correct Poisson subsampling assumed by the privacy accounting methods, and therefore the implementation might have significantly weaker privacy properties than claimed (Lebeda et al., 2024).
To make a fair comparison between all methods, we implement the Poisson subsampling, the same way Opacus does it, for all frameworks and a new custom Batch Memory Manager to flag when it is time to take a step. This way, all the experiments are seeded and have the same logical batch sizes.
2.2 DP-SGD Gradient Clipping Optimizations
@@
Ghost clipping computes the loss gradient norm after the backpropagation optimization and then reweights the loss to update the clipped gradients. Its main contribution is memory saving at the cost of adding another backward pass (Li et al., 2022).
Mixed Ghost clipping (Bu et al., 2022) builds on-top of Ghost clipping. It implements the ghost clipping technique for convolutional layers. Its main contribution is that the algorithm will decide when to clip the gradients using per-example or ghost. This difference matters because the ghost clipping is less efficient when the layer’s input dimensions are too big. E.g., for ResNets, each clipping method will be applied for half of the layers. The first layers will be clipped using the the per-example and then ghost clipping in the bottom layers. As the model goes deeper, the feature size decreases, and the number of channels increases, prioritizing ghost clipping (Bu et al., 2022).
Book Keeping (Bu et al., 2023) uses all the previous techniques but requires only one backpropagation pass without explicitly calculating the per-example gradients. It avoids the second pass that ghost clipping does by reusing the intermediate results of the output gradients to calculate the sum of the clipped gradients and the clipping factor. Book Keeping can also be implemented together with the mix optimization. It also implements another technique called MixOpt, which does the same evaluation as the mix ghost clipping, but also determines whether doing a second backward pass is more efficient.
Implementations of optimized DP-SGD PrivateVision (Bu et al., 2022), and FastDP (Bu et al., 2023) are PyTorch-based implementations. While both PrivateVision and FastDP implement ghost clipping and its variants, all FastDP implementations use the Book Keeping method. By the time of the submission, Opacus ghost clipping implementation was not ready for production. We tested it but excluded it from the analysis.
JAX We compare all implementations with two JAX implementations: The first one is a native JAX (Bradbury et al., 2018) implementation that clips the per-sample gradients with Optax (DeepMind et al., 2020) without utilizing any further optimization. This naive implementation in JAX is prone to recompiling due to changing tensor sizes caused by the Poisson subsampling. The second implementation which we call Masked DP-SGD overcomes this issue by processing always full physical batches and then masking gradients of samples that are not required (see Algorithm 2). This implementation computes always slightly more gradients but is not prone to repeated recompilation.
@@ 3 Experiment Overview
Metrics We compare the throughput, defined as how many samples can be processed per second during training, and the maximum physical batch size, reached before running out of memory.
Dataset We benchmark with the CIFAR100 (Krizhevsky & Hinton, 2009) resized to 224x224.
Models We train two families of models: Vision Transformer (ViT) (Dosovitskiy et al., 2021) and ResNet (Kolesnikov et al., 2020) (See Table 1). Both are pre-trained on ImageNet-21k (Russakovsky et al., 2015).
| Vision Transformer (ViT) | ResNet | ||
| Type | # of Parameters | Type | # of Parameters |
| Tiny | 5.7 M | 23.7 M | |
| Small | 22.1 M | 42.7 M | |
| Base | 86.6 M | 211.8 M | |
| Large | 304.3 M | 382.4 M | |
| Huge | 630.8 M | 929.2 M | |
Parameterization While parameter-efficient fine-tuning of some parts of the model has been shown to be effective under DP (Tobaben et al., 2023; Yu et al., 2022), our work focuses on the computational efficiency of DP-SGD and thus we consider the worst-case scenario of fine-tuning all parameters of the model. Furthermore, any training from scratch requires training all parameters.
Hyperparameters We train for four optimizer steps with a sampling rate of (expected batch size of 25000), which allows us to test the experiments quickly with a realistic high batch size (Ponomareva et al., 2023; Räisä et al., 2024). We do not focus on finding the best possible utility, which requires training for many more epochs (See Table A3 for the accuracy after training for four steps).
Environment specifications We use two GPU architectures: NVIDIA V100 (32 GB VRAM) and A100 (40 GB VRAM) with identical Python environments. Each node contains four GPUs. We use 16 CPU workers for data loading. In the distributed case of more than one GPU, we cannot use multiple workers.
Source code We provide the code for reproducing the experiments at https://github.com/DPBayes/Towards-Efficient-Scalable-Training-DP-DL.
@@ 4 What is the cost of DP in Deep Learning?
In this section we will quantify the computational cost of deploying DP training. We do this by comparing the throughputs and maximum physical batch sizes between the non-private training with PyTorch and private training with Opacus, which is the most used DP-SGD implementation. Additionally, we identify the reasons for the higher computational cost of DP-SGD through profiling.
4.1 Throughput and Maximum Batch Size Comparison
We compare relative throughput (Figure 2) and the maximum physical batch size (Figure 3) between DP-SGD (Opacus) and non-private training with PyTorch. The main metric of interest is the throughput as it quantifies the training speed but the maximum physical batch size becomes important when training models that are too large to fit even one example at a time. For both metrics DP-SGD becomes more expensive with larger models but the detailed trends differ.
@@ 4.1 Throughput and Maximum Batch Size Comparison
Vision Transformer The throughput difference between Opacus and the non-private baseline with PyTorch (see Figure 2(a)) does not spike but grows steadily as a function of model size, which is interesting considering how big the relative difference is in maximum physical batch size (Figure 3(a)) is: The throughput ranges from a relative difference of for the smallest model to for the largest model while the maximum physical batch size has a relative difference of around for the smallest model and for the largest model.
@@
ResNets We observe a less steady growth trend in terms of throughput with the ResNets in Figure 2(b). When increasing the model size the ResNets do have spikes of growth as the model size grows. The contrast in Figure 2 between ViT and BiT ResNet models is due to the architecture and types of layers. The parameter space grows as the width factor (see Table 1) for the ResNets, so the makes the neural network wider by a factor of three. Based on our results, the width of the layers affects throughput much more than the depth of the network. They have comparable throughput with the same width and different depths, but increasing the width will make the model in the private setting much slower and reduce the maximum batch size significantly.
@@
How much does finding the maximum physical batch size matter? In Figure A.1 in the Appendix, we display the relative throughput as a percentage by dividing the throughput at a particular physical batch size by the maximum achievable throughput. We see that as the physical batch size increases, the throughput will grow as expected, but at some point there is no significant further improvement in throughput from using a larger physical batch size. Practitioners may estimate the optimal batch size based on available memory and performance trade-offs. It is not crucial to set the physical batch size to the maximum possible but a good enough value is fine. Typically, throughput using smaller batches is limited by data loading speeds, but as batch size increases, computation becomes the limiting factor.
@@ 4.2 Reasons for the Increase in Computational Cost
Giving a detailed breakdown of low-level operations associated with DP is challenging. However, using GPU profiling tool NVIDIA Nsight System, we can identify three aspects which constitute the majority of DP overheads. Firstly, due to its larger memory footprint, DP-SGD is able to consume smaller physical batches than its non-private counterpart. This results in larger amount of smaller low-level kernel calls which leads to slightly lower utilisation of the GPU. At very small batch sizes even the kernel launch overheads can become a notable factor for slowdown. Secondly, the computation of per-example gradients introduces significant overhead in the backward pass as it cannot be parallelised as in batched gradient computation. This is the most prominent cause of the total overhead. Finally, an additional DP-optimizer step that clips and accumulates the per example gradients, which is not present in the non-DP algorithm, needs to taken after each physical batch (see Table 2).
| Section | PyTorch non-private | Opacus per-example |
| Forward | 81.14 | 101.53 |
| Backward | 163.85 | 681.48 |
| Clip and Accumulation | 0 | 26.76 |
| Optimizer Step | 38.17 | 99.65 |
@@ 5 Decreasing the computational cost
5.1 Optimized gradient clipping algorithms
First, we evaluate the more efficient gradient clipping implementations that have been described in Section 2.2 using the Vision Transformer ViT base model. We chose it as our benchmark model because the middle model size is large enough to evaluate the advantages of the optimized gradient clipping algorithms but does not require excessive amount of time to train. The non-Opacus implementations do not support the BiT ResNet due to their custom weight standardization layer.
@@ 5.1 Optimized gradient clipping algorithms
Throughput Comparison Figure 4 displays the throughput for each clipping algorithm for each tested GPU. Moving from a V100 to an A100 GPU increased the throughput by times on average over all clipping methods. The one that benefited the most is the per-example clipping by Opacus with a improvement in throughput. This is because of Opacus-specific optimizations. Their implementation is optimized to vectorize the virtual batches and get the most out of the processing unit to compensate for the per-example clipping. Also, we base our virtual batching module on Opacus, which may have further contributed to the advantage seen for Opacus. The other implementations showed benefits similar to those of non-private training. In both GPU architectures, the clipping optimizations consistently maintained their relative throughput difference to their non private baseline. Private Vision gets closer to the non-private baseline physical batch size, but Book Keeping is closer to its throughput with a smaller physical batch size (see Figure 6).
@@
5.2 Lower precision
We consider using lower precision to speed up computation. We evaluate the use of Tensor Core 32 (TF32) for training. TF32 has 10 bits for precision, with eight range bits, giving it the same range but less precision than 32-bit single precision floats (FP32) (Kharya, 2020). Using lower precision can have benefits exactly where DP training struggles: it requires less memory, uses fewer bits to represent the data, and its operations are optimized for GPU, making them much faster (NVIDIA, 2023). It is specially optimized for the A100 GPU and unavailable for the V100, so we compared training on the A100 with and without TF32.
Experimental results In Figure 5, we display the relative difference between mean throughput between runs with TF32 and FP32. For non-private training, throughput increases with model size. For private training throughput increases for the smaller models, but it goes down again as the model size grows after the base size. Models that are too small do not gain much from TF32, and the larger ones are too expensive. Regarding the memory advantages by TF32, we could not see an improvement. Both models, with and without TF32, could fit the same number of instances.
Concerns regarding TF32 under DP There are two concerns with using lower precision in DP deep learning: its effects on utility and privacy. For the first issue, using lower precision may affect utility, as it is less precise. We did not find a significant decay in the accuracy of the models; it differs by decimal points at the precision (See Table A3 in the Appendix). Regarding privacy, all floating point implementations provide imperfect implementations of real-valued mechanisms, and this can cause additional privacy vulnerabilities (Mironov, 2012). Using lower precision may exacerbate the problem. Discrete mechanisms (e.g. Canonne et al., 2020; Agarwal et al., 2021) that avoid the theoretical challenges exist, but they are often less convenient and may lead to loss of utility, especially in low precision settings. The efficiency of different discrete mechanisms in TF32 is an interesting topic of further research.
6 JAX